# Towards Safe and Robust Autonomous Vehicle Platooning: A Self-Organizing Cooperative Control Framework

### v1

@@ Abstract
In the emerging hybrid traffic flow environment, which includes both human-driven vehicles (HDVs) and autonomous vehicles (AVs), ensuring safe and robust decision-making and control is crucial for the effective operation of autonomous vehicle platooning. Current systems for cooperative adaptive cruise control and lane changing are inadequate in responding to real-world emergency situations, limiting the potential of autonomous vehicle platooning technology. To address the aforementioned challenges, we propose a Twin-World Safety-Enhanced Data-Model-Knowledge Hybrid-Driven autonomous vehicle platooning Cooperative Control Framework. Within this framework, a deep reinforcement learning formation decision model integrating traffic priors is designed, and a twin-world deduction model based on safety priority judgment is proposed. Subsequently, an optimal control-based multi-scenario decision-control right adaptive switching mechanism is designed to achieve adaptive switching between data-driven and model-driven methods. Through simulation experiments and hardware-in-loop tests, our algorithm has demonstrated excellent performance in terms of safety, robustness, and flexibility. A detailed account of the validation results for the model can be found in Our Website.
@@
Autonomous Vehicle Platooning; Deep Reinforcement Learning; Mixed Traffic Flows; Self-Organised Cooperative Control.
@@ I Introduction
The advent of autonomous driving and vehicle communication technologies has sparked significant interest in connected self-driving vehicles. Among these, autonomous vehicle platooning stands out for its potential to enhance traffic efficiency and safety [1, 2] in a relatively straightforward and cost-effective manner. By reducing aerodynamic drag, vehicle platooning can lower fuel consumption and emissions by at least 10%[3, 4],thereby alleviating environmental pollution[5, 6], while also decreasing road space usage and traffic congestion[7].
Most existing studies on autonomous vehicle platooning primarily focus on cooperative adaptive cruise control (CACC) and the cooperative lane change control (CLCC) to achieve a narrower inter-vehicle gap and a lower fuel consumption. Current autonomous driving platoons generally use LiDAR for virtual connectivity and communicate via wireless communication technologies[8]. Different communication protocols, including DSCR[9], GPSR[10], and GAHG[11], are utilized in various platooning algorithms. Different platoon structures have also been explored, which can be categorized into fully autonomous platoons and heterogeneous platoons with partially HDVs. The heterogeneous platoons can be further divided into those where the lead vehicle is human-driven and those where both the lead and tail vehicles are human-driven. Given the future development trends, this study primarily investigates control strategies for fully autonomous vehicle platoons.
CACC systems are integral for maintaining reduced inter-vehicle gaps and achieving lower fuel consumption, with a primary emphasis on ensuring the string stability[12, 13] of the platoon. These systems can be classified into model-based and rule-based methods. Model-based approaches, such as Model Predictive Control (MPC)[14, 15, 16, 17], leverage vehicle dynamics models and optimization techniques to achieve precise control of acceleration and deceleration[18]. However, they are often computationally demanding and may not adapt effectively to rapidly changing conditions. In contrast, rule-based methods rely on predefined rules and logic to dictate vehicle behavior[19, 20, 21], offering simplicity and faster computation at the cost of reduced flexibility in complex environments. In addition to these traditional methods, data-driven approaches, particularly those based on Reinforcement Learning (RL), have emerged as a promising alternative[13]. RL-based CACC systems can optimize control policies by learning from real-time data[22], effectively managing complex dynamics and improving response to changing traffic conditions. Despite their potential, these methods require substantial training data and pose challenges in ensuring safety during the learning phase.
CLCC enhances the maneuverability and efficiency of vehicle platoons by enabling coordinated lane changes. Advanced CLCC strategies, similar to CACC, include both model-based and rule-based approaches[23, 24, 25]. Model-based methods involve predictive control techniques to optimize lane change maneuvers based on anticipated future vehicle states[26, 27], while rule-based methods utilize a set of established rules to manage lane changes according to current traffic conditions[28, 29, 30]. Nevertheless, both approaches face limitations; model-based methods can be computationally intensive and sensitive to model inaccuracies, while rule-based methods may lack the adaptability required for managing complex traffic scenarios involving multiple interacting vehicles. RL-based approaches in CLCC similarly offer greater adaptability by learning from diverse traffic scenarios, allowing for optimized lane-changing strategies[31]. While these methods enhance flexibility, they require significant data and computational resources, and ensuring reliability in mixed traffic environments remains a challenge.
Current CACC and CLCC methods have notable shortcomings, particularly in handling scenarios with multiple types of strong interference. Existing strategies are typically designed to cope with one or two types of disturbances, such as sensor noise or communication delays, but struggle when faced with more complex interference patterns [32]. This limitation can lead to safety-critical situations, especially in mixed traffic environments where unpredictable HDVs are present. Consequently, the current state of autonomous vehicle platooning remains relatively conservative, often prioritizing safety over efficiency to mitigate risks.
To address these limitations, insights can be drawn from the robotics and Unmanned Aerial Vehicle (UAV) domain, where emergency obstacle avoidance and platoon reconfiguration strategies have been actively researched [33, 34, 35]. In robotics, these strategies often employ techniques such as real-time path planning[36, 37], dynamic reconfiguration algorithms, and machine learning-based decision-making to ensure robust performance in uncertain environments. These methods have demonstrated success in controlled and relatively closed environments, such as factory floors or isolated test tracks.
However, the application of these advanced strategies to autonomous vehicle platoons is limited by the open and highly variable nature of mixed traffic environments. Unlike controlled environments, mixed traffic flows are subject to unpredictable human behavior and diverse vehicle types, which impose significant challenges on the direct transfer of robotic strategies to autonomous driving. Despite these challenges, adapting and refining these approaches for use in vehicle platooning could potentially enhance the robustness and safety of autonomous driving systems.
In light of the aforementioned challenges, this study proposes a novel approach to autonomous vehicle platooning control. We propose a Twin-World Safety-Enhanced Data-Model-Knowledge Hybrid-Driven Adaptive Autonomous Vehicle Platooning Cooperative Control Framework. This framework achieves adaptive decision-making for platooning in various complex scenarios by flexibly switching between deep reinforcement learning (DRL) integrated with traffic prior knowledge and model control methods. Firstly, the vehicle platooning problem is modeled as a multi-agent cooperative decision-making problem, and a RL decision-making framework based on Proximal Policy Optimization (PPO)[36] is proposed, with a decision network structure and reward function designed based on traffic prior knowledge. A twin-world model based on safety priority judgment is introduced and coupled with the DRL algorithm to enhance algorithm safety. Subsequently, combining optimal control methods, a multi-scenario decision-control right adaptive switching mechanism is designed. To evaluate our method, we have devised intricate and high-risk mixed-traffic scenarios, as depicted in Figure 1. These scenarios encompass various elements such as human interference, traffic accidents, and fluctuations in traffic flow. Through simulations and hardware-in-the-loop experiments, our algorithm has been validated in terms of safety, robustness, and flexibility.
Figure 1: Illustration of the considered traffic scenario and the solution. CAVs (Red) and HDVs (Green) are both considered
Our contributions are summarized as follows:
  • 1)
@@
    A Data-Model-Knowledeg Hybrid-Driven Framework for autonomous vehicle platooning is constructed, which integrates data-driven methods, model-based control, and domain-specific knowledge, enabling autonomous vehicle platoons to achieve adaptive decision-making and robust control in complex and dynamic traffic environments.
  • 2)
    A twin-world safety enhancement model is coupled with RL and implement an adaptive switching strategy. The mechanism enables the platoon to seamlessly switch between routine and safety-critical operations.
  • 3)
    A Cyber-Physical Experimentation Platform is established, incorporating simulation and real-world controllers, to validate the framework’s robustness and feasibility under diverse environmental conditions.
@@
To better present our work, the rest of this paper is organized as follows. Section II presents the hybrid-driven self-organizing control framework. Section III introduces in detail the proposed Twin-World Safety-Enhanced Data-Model-Knowledge Hybrid-Driven Adaptive Autonomous Vehicle Platooning Cooperative Control Framework Section IV presents the performance evaluation. Finally, we conclude the paper in Section V.
II Problem statement and framework review
@@ II-A Scenario Description
As the most common application scenario for autonomous vehicle platooning, we have conducted an in-depth study on the self-organizing decision-making problem of multi-vehicle platooning on a three-lane highway with strong disturbances. Based on real-world issues encountered in various platooning tests, we identified and summarized several major types of strong disturbances, including human interference, traffic accidents, and traffic flow oscillations. In order to address these disturbances, a series of representative scenarios were devised, including emergency braking of the leading vehicle, sudden traffic accidents, and medium to low-speed traffic flow scenarios. These are illustrated in Figure 1.
Conventional obstacle avoidance strategies for autonomous vehicle platooning often struggle to handle these strong disturbance factors. For the practical implementation of autonomous vehicle platooning, it is essential to safely navigate through regions with strong disturbances and ultimately maintain stable operation at high speeds to reach the destination.
In this study, the high-level action decisions for the autonomous vehicle platoon are determined by our proposed Twin-World Safety-Enhanced Data-Model-Knowledge Hybrid-Driven framework. This algorithm then utilizes a closed-loop Proportional-Integral-Derivative (PID) controller to translate these decisions into low-level steering and acceleration/deceleration controls.
II-B Framework Overview
Figure 2: The overview of the proposed data-model-knowledge hybrid-driven framework for self-organizing autonomous vehicle platooning
To address emergency risk scenarios that traditional vehicle platooning technologies cannot handle, this study proposes a Twin-World Safety-Enhanced Data-Model-Knowledge Hybrid-Driven Adaptive Autonomous Vehicle Platooning Cooperative Control Framework. As illustrated in Figure 2, the framework is structured into two primary layers:
  • Upper Layer: This layer utilizes a Proximal Policy Optimization DRL decision model following the Centralized Training and Decentralized Execution (CTDE) paradigm. The DRL model is coupled with a twin-world model to perform safety verifications and ensure the generation of safe decisions.
  • Lower Layer:This layer implements a Linear-Quadratic Regulator (LQR) for the longitudinal control of platoon vehicles, aimed at maintaining smaller inter-vehicle headway distances.
These decision-making strategies are adaptively and flexibly switched using a knowledge-based finite state machine. The LQR controller is employed to maintain smaller headway distances in safe driving environments, thereby enhancing the platoon’s efficiency. In complex and challenging scenarios, the reinforcement learning model assumes control, managing the splitting and reformation of the platoon to ensure more flexible and safe driving.
@@ III Methodology
This section details the algorithm design for autonomous vehicle platoon reconfiguration and maintenance. First, the problem is modeled as a Partially Observable Markov Decision Process (POMDP)[38], with a clear explanation of the observation space, action space, and reward function, integrating prior traffic knowledge. Then, the safety-priority-based twin-world model is introduced. Finally, this study combines optimal control methods to design an adaptive switching mechanism suited for multi-scenario applications, referred to as the platoon maintenance model.
III-A Problem Formulation and Reinforcement Learning Setup
Driving platoon vehicles on highways is a complex and challenging task, especially in mixed traffic flows. Treating the platoon as a single entity can lead to safety risks under strong disturbances, requiring reconfiguration. This process involves multi-agent cooperative decision-making[39], as each vehicle’s state is influenced by the actions of other platoon members and surrounding traffic.
Traditional decision algorithms are inadequate for this task; thus, the problem should be modeled as a POMDP to employ deep reinforcement learning. The challenge lies in balancing individual and collective goals while navigating a dynamic environment where actions influence both nearby and distant vehicles. The system’s schematics are shown in Figure 3.
In this study, we model the problem of multi-vehicle platooning on highways as an agent that accesses an observation at each time $t$. Given the sensor capabilities of platoon vehicles and the complexity of real traffic environments, the observation only provides partial information about the system state $s_{t}$. The platoon cannot accurately perceive the actions and intentions of all other vehicles, making the observation space $o_{t}$ a proper subset of the system state $s_{t}$.
Using $o_{t}$, the platoon can sample decision actions $a_{t}$ from the action space according to the shared decision policy $\pi_{\theta}$, as follows:
$a_{t}\sim\pi_{\theta}\left(a_{t}\mid o_{t}\right)$ (1)
where the variable $\theta$ represents the trained policy parameters. The agents within the platoon execute action $a_{t}$, interact with the environment, and transition the system to state $s_{t+1}$, acquiring a new observation $o_{t+1}$.
The vehicle platoon, while making decisions as a whole, employs a decentralized execution strategy. Therefore, we can formulate this problem as a POMDP, which can be described by the tuple $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\mathcal{O})$, where $\mathcal{S}$ is the set of system states, $\mathcal{A}$ is the action space, $\mathcal{P}$ is the state transition model, $\mathcal{R}$ is the reward function, and $\mathcal{O}$ is the observation space. The observation space $\mathcal{O}$ is determined by the system state and the position of the platoon $({o_{t}}\in{s_{t}})$.
In the POMDP, the agent makes decisions based on the policy $\pi$ and the current observation space. The objective of the agent is to find the optimal policy $\pi^{*}$ that maximizes the cumulative future rewards[40]. The optimal action-value function $Q^{\pi}(s,a)$, which can be defined as follows:
The subsequent section will present the design of the observation space $\mathcal{O}$, action space $\mathcal{A}$, and reward function $\mathcal{R}$.
  • Action space: In this study, each agent $\text{agent}_{i}$ has an action space that includes a set of high-level control decisions: left turn, right turn, accelerate, decelerate, and cruise[41, 42]. For vehicle platooning, we integrate the action spaces of multiple agents using a quinary encoding scheme. This approach encodes the combined action space using a shared encoder, forming a 125-dimensional discrete action space through the Cartesian product of the individual actions. The overall action space for the platoon is thus the joint action space of all autonomous vehicles, denoted as $\mathcal{A}=\mathcal{A}_{1}\times\mathcal{A}_{2}\times\cdots\times\mathcal{A}_{N}$.
    Figure 3: The schematics of the system and the simulation setup
    As illustrated in Figure 3, after making high-level decisions, the platoon’s bottom-level controllers manage each vehicle’s steering and throttle, generating specific control signals to execute vehicle movements in the simulation environment.
  • Observation Space: The observation space can be represented as a matrix with dimensions $|N_{i}|\times|F|$, where $|N_{i}|$ is the number of vehicles observed by the agent, and $|F|$ represents the feature vector of other vehicles’ positions and actions. The feature vector can be defined as:
    $F_{i}=[x_{i},y_{i},v_{i}^{x},v_{i}^{y},is\_in\_platoon]$ (3)
    To better represent the interactions between the platoon and surrounding vehicles, $x_{i},y_{i},v_{i}^{x},v_{i}^{y}$ are the relative positions and velocities of other vehicles with respect to the observing vehicle, indicating longitudinal position, lateral position, longitudinal velocity, and lateral velocity, respectively. The variable ${is\_in\_platoon}$ is a binary indicator (0 or 1) that denotes whether the observed vehicle is part of the platoon.
    In this study, we consider the sensor detection range limitations and assume that only nearby vehicles can be observed. Specifically, only vehicles within $D_{vision}$ ahead or behind the platoon can be detected[43]. The observation space for the entire platoon system is the Cartesian product of the individual vehicles’ observation spaces, denoted as $\mathcal{S}=\mathcal{S}_{1}\times\mathcal{S}_{2}\times\cdots\times\mathcal{S}_{N}$.
  • Reward Function: The reward function is a crucial component in training the DRL model, as it determines the model’s overall performance. This study designs both single-vehicle rewards and multi-vehicle collaborative rewards to ensure the robustness and safety of the platoon’s decisions. The overall reward for the vehicle platoon is the sum of the rewards obtained by each vehicle and the collective rewards:
    $R_{global}(s,a)=\frac{r_{L}}{n}\sum_{i=1}^{n}R_{ind}^{i}+R_{sys}$ (4)

### v2

# Towards Safe and Robust Autonomous Vehicle Platooning: A Self-Organizing Cooperative Decision-making Framework

@@ Abstract
In hybrid traffic environments where human-driven vehicles (HDVs) and autonomous vehicles (AVs) coexist, achieving safe and robust decision-making for AV platooning remains a complex challenge. Existing platooning systems often struggle with dynamic formation management and adaptability, especially in unpredictable, mixed-traffic conditions. To enhance autonomous vehicle platooning within these hybrid environments, this paper presents TriCoD, a twin-world safety-enhanced Data-Model-Knowledge Triple-Driven Cooperative Decision-making Framework. This framework integrates deep reinforcement learning (DRL) with model-driven approaches, enabling dynamic formation dissolution and reconfiguration through a safety-prioritized twin-world deduction mechanism. The DRL component augments traditional model-driven methods, enhancing both safety and operational efficiency, especially under emergency conditions. Additionally, an adaptive switching mechanism allows the system to seamlessly shift between data-driven and model-driven strategies based on real-time traffic demands, thereby optimizing decision-making ability and adaptability. Simulation experiments and hardware-in-the-loop tests demonstrate that the proposed framework significantly improves safety, robustness, and flexibility. A detailed account of the validation results for the model can be found in Our Website.
@@
Autonomous Vehicle Platooning; Deep Reinforcement Learning; Mixed Traffic Flows; Self-Organised Cooperative Decision-making.
@@ I Introduction
The rapid advancements in autonomous driving and vehicle communication technologies have catalyzed significant interest in connected and autonomous vehicles (CAVs). Among these technological advances, autonomous vehicle platooning has emerged as a promising approach to enhance traffic efficiency, safety, and environmental sustainability [1, 2]. By reducing inter-vehicle gaps and aerodynamic drag, vehicle platooning holds substantial potential to decrease fuel consumption and emissions[3], addressing environmental challenges and mitigating traffic congestion[4, 5].
Current autonomous vehicle platooning systems primarily focus on two key aspects of vehicle control: Cooperative Adaptive Cruise Control (CACC) and Cooperative Lane Change Control (CLCC). CACC manages longitudinal dynamics to maintain inter-vehicle gaps and ensure string stability [6, 7], while CLCC facilitates lateral maneuvers for lane changes within the platoon. Extensive research has been dedicated to improving CACC and CLCC. Model-based approaches, such as Model Predictive Control (MPC), have been employed to optimize vehicle dynamics [8, 9, 10]. However, these approaches often struggle with computational complexity and adaptability to rapidly changing environments. Rule-based methods offer simplicity but lack flexibility [11, 12, 13], while Reinforcement Learning (RL) has recently emerged as a promising alternative capable of learning from real-time traffic data [7, 14].
Despite these contributions, CACC and CLCC face intrinsic limitations in managing dynamic interactions and formation reconfiguration within mixed traffic environments[15, 16], restricting their applicability to real-world scenarios characterized by complex and unpredictable conditions [17]. Addressing these challenges necessitates a more dynamic and adaptable approach to formation maintenance and reconfiguration.
To overcome these challenges, inspiration is drawn from formation decision-making techniques in other domains, specifically Unmanned Aerial Vehicles (UAVs). In the broader context of multi-agent systems, the concept of formation reconfiguration has been extensively explored in UAVs [18], which leverage dynamic formation reconfiguration to adapt to changing conditions, avoid obstacles, and fulfill mission requirements in controlled three-dimensional environments[19, 20]. Although UAVs operate in different environments compared to ground vehicles, certain principles, such as decentralized control and dynamic adaptability, can be applied to autonomous vehicle platoons to improve their formation decision capabilities.
Unlike UAVs, which operate in three dimensions with relatively low obstacle densities, autonomous vehicle platoons must navigate constrained road environments with frequent interactions involving human-driven vehicles (HDVs) and complex infrastructure[21]. The unpredictable behavior of HDVs, combined with road constraints, presents a significant challenge for dynamic formation reconfiguration in vehicle platoons[22]. Addressing these challenges requires a novel approach that integrates adaptability with the specific needs of ground-based traffic systems.
In response, this paper proposes TriCoD, a safety-enhanced Data-Model-Knowledge Triple-Driven Adaptive Cooperative Decision-making Framework for autonomous vehicle platooning, integrating data-driven methods with model-based control to manage platoons in complex environments. The platooning problem is modeled as a multi-agent cooperative decision-making challenge based on Trust Region Policy Optimization (TRPO) [23], with a twin-world safety enhancement model runs in parallel with real scenarios to improve safety and adaptability during learning. An adaptive switching mechanism optimizes transitions between decision strategies based on traffic conditions.
The hybrid approach is designed to enable platoon reconfiguration in complex traffic conditions. The system dynamically adjusts both longitudinal and lateral behaviors to enhance safety, efficiency, and flexibility in real time. The approach is evaluated in high-risk mixed-traffic scenarios, including human interference, incidents, and fluctuating flows, as shown in Figure 1. Extensive software-in-the-loop and hardware-in-the-loop experiments fully demonstrate that our framework exhibits higher security and robustness in different scenarios while taking efficiency into consideration.
The contributions of this paper are summarized as follows:
@@
  • A hybrid framework for platooning decision has been developed, which integrates data-driven methods with model-based control and domain-specific knowledge, enabling platoon members to achieve adaptive decision-making in complex traffic environments.
  • A twin-world safety enhancement model integrated with RL is proposed, which significantly enhances the sampling efficiency by providing a parallel virtual environment that allows safe exploration of diverse strategies, thereby improving the overall safety and adaptability of the system.
  • The proposed framework uniquely enables the dissolution and reconfiguration of platoon formations in response to changing traffic conditions. It integrates longitudinal and lateral control, allowing the system to dynamically adjust vehicle behaviors and maintain formation integrity in complex environments.
@@
To better present our work, the rest of this paper is organized as follows. Section II formulates the problem in our research. Section III introduces in detail the proposed TriCoD framework. Section IV presents the performance evaluation. Finally, the paper is concluded in Section V.
Figure 1: Illustration of challenging traffic conditions and our framework’s adaptive responses: formation maintenance, dissolution, and reorganization.CAVs (red) and HDVs (green) are represented with corresponding behaviors under each scenario.
II Problem statement
@@ II-A Scenario Description
Highway environments are the most common application scenario for autonomous vehicle platooning. This study focuses on self-organizing decision-making for multi-vehicle platooning on a three-lane highway, addressing significant disturbances such as human interference, traffic accidents, and traffic flow oscillations, as illustrated in Figure 1. In these scenarios, traditional formation algorithms are no longer capable of effectively addressing the challenges posed by unpredictable disturbances, leading to reduced stability and compromised safety. Therefore, a new, adaptive algorithm is necessary to navigate these disturbances and ensure robust, stable, high-speed operation.
II-B Problem Formulation
The problem is modeled as a Partially Observable Markov Decision Process (POMDP)[24], which provides a robust framework for handling the partial observability and stochastic nature of real-world traffic scenarios. By employing DRL within the POMDP framework, both individual and collective goals can be balanced, enabling adaptive responses in dynamic environments. The system’s schematics are shown in Figure 2.
The vehicle platoon, while making decisions as a whole, employs a decentralized execution strategy. The problem is formulated as a POMDP, which can be described by the tuple $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\mathcal{O})$, where $\mathcal{S}$ is the set of system states; $\mathcal{A}$ is the action space; $\mathcal{P}$ is the state transition model; $\mathcal{R}$ is the reward function; and $\mathcal{O}$ is the observation space.
In this POMDP, the agent makes decisions based on the policy $\pi$ and the current observation space. The optimization problem is then formulated as finding the optimal policy $\pi^{*}$ that maximizes the expected sum of discounted future rewards[25]:
where $\gamma\in[0,1]$ is the discount factor that balances the importance of immediate versus future rewards. The optimization process involves updating the policy $\pi_{\theta}$ to approximate the optimal policy $\pi^{*}$, which governs how the platoon should act to achieve long-term performance goals.
II-B1 Observation Space $\mathcal{O}$
In real-world highway environments, autonomous vehicles cannot directly observe the full system state due to sensor limitations and occlusions caused by other vehicles. Formally, the observation $o_{t}$ at time $t$ is a function of the true state $s_{t}$, but is incomplete:
$o_{t}=f(s_{t}),\quad o_{t}\subset s_{t}$ (2)
In this study, the observation space is structured as a matrix with dimensions $|N_{i}|\times|F|$, where $|N_{i}|$ is the number of vehicles detected by the platoon, and $|F|$ represents the feature vector describing each detected vehicle. The feature vector includes relative positions and velocities of other vehicles, as well as an indicator variable $\delta_{i}^{platoon}$ that specifies whether the vehicle is part of the platoon:
$F_{i}=[x_{i},y_{i},v_{i}^{x},v_{i}^{y},\delta_{i}^{platoon}]$ (3)
where $x_{i}$ and $y_{i}$ represent the relative longitudinal and lateral positions of vehicle $i$, and $v_{i}^{x}$ and $v_{i}^{y}$ denote its relative longitudinal and lateral velocities. The binary variable $\delta_{i}^{platoon}$ indicates whether vehicle $i$ is part of the platoon:
$\delta_{i}^{platoon}=\begin{cases}1&\text{if vehicle }i\text{ belongs to the platoon},\\ 0&\text{otherwise}.\end{cases}$ (4)
Additionally, the observation is restricted to vehicles within a sensor detection range, $D_{vision}$, ahead or behind the observing vehicle.
Figure 2: Illustration of the system and the simulation setup. All high-level actions are verified through a security mask before being processed by the bottom-level controller, which will be finally converted into control signals.
II-B2 Action Space $\mathcal{A}$
The action space $\mathcal{A}$ defines the set of possible actions that the platoon can take at any given time step. In this study, each autonomous vehicle in the platoon selects from a discrete set of high-level actions, including left turn, right turn, accelerate, decelerate, and cruise. These actions allow the vehicles to maneuver within their lanes and adjust their speeds to maintain a safe and efficient formation.
For a platoon consisting of $N$ autonomous vehicles, the joint action space $\mathcal{A}$ is the Cartesian product of the individual action spaces of the vehicles:
$\mathcal{A}=\mathcal{A}_{1}\times\mathcal{A}_{2}\times\cdots\times\mathcal{A}_{N}$ (5)
Each vehicle $i$ in the platoon selects an action $a_{i}\in\mathcal{A}_{i}$, and the combined action $a_{platoon}$ is executed as:
$a_{platoon}=(a_{1},a_{2},\ldots,a_{N})$ (6)
In this study, a quinary encoding scheme is employed to encode the joint action space, resulting in a 125-dimensional discrete action space through the Cartesian product of the individual actions. After selecting high-level actions, the platoon’s low-level controllers manage steering and throttle to execute the specific vehicle movements in the simulation environment, as illustrated in Figure 2.
II-B3 Transition Probabilities $\mathcal{P}$
HDVs introduce randomness into the whole system, making future states inherently uncertain. This uncertainty is modeled using a state transition function $\mathcal{P}(s_{t+1}|s_{t},a_{t})$, which represents the stochastic nature of the environment, capturing the likelihood of transitioning from the current state $s_{t}$ to the next state $s_{t+1}$ given the current action $a_{t}$:
$\mathcal{P}(s_{t+1}|s_{t},a_{t})=\Pr(s_{t+1}|s_{t},a_{t})$ (7)
The transition probabilities account for the stochasticity brought by the HDVs, ensuring that the decision-making process anticipates and responds to these uncertainties.
@@ III Methodology
This section introduces a comprehensive framework for autonomous vehicle platoon. The DRL-based core decision-making algorithm for formation reconfiguration is first presented to enable adaptive and secure platoon adjustments, which is followed by a formation maintenance algorithm and an adaptive switching mechanism, ensuring stable and efficient decision-making across varying traffic conditions.
III-A Framework Overview
Figure 3: Overview of the TriCoD framework for self-organizing autonomous vehicle platooning, which combines a data-driven upper layer and a model-driven lower layer to manage platoon adaptively. The upper layer uses DRL with a twin-world safety verification model for safe, real-time decision-making, while the lower layer employs a LQR for precise control in predictable conditions, ensuring stable and efficient vehicle spacing. Together, these layers provide robust, adaptive platooning in complex traffic environments.
To address emergency and high-risk scenarios that traditional vehicle platooning technologies struggle to manage effectively, this study introduces TriCoD framework. As depicted in Figure 3, the framework seamlessly integrates two tightly coupled primary layers that collaboratively handle different aspects of platoon decision making:
  • Upper Layer: The DRL decision model follows the Centralized Training and Decentralized Execution (CTDE) paradigm, with a twin-world safety verification model that simulates decision outcomes in a parallel virtual environment. This verification ensures safe, effective decision-making before deployment and enhances the DRL model’s sampling efficiency. The upper layer works alongside the lower layer to adapt to complex scenarios in real-time.
  • Lower Layer: The Linear-Quadratic Regulator (LQR) provides precise control in predictable scenarios, minimizing headway for efficiency. This layer also collaborates with the DRL model by executing low-level control commands that align with the strategic decisions made by the upper layer.
A knowledge-based finite state machine dynamically integrates these layers, switching decision strategies based on real-time traffic conditions. In stable conditions, the LQR controller optimizes efficiency with reduced headway. In complex scenarios, the DRL model manages platoon splitting and reformation, coordinating with the LQR for adaptive responses. This hybrid approach combines data-driven learning with model-based control, ensuring cohesive, adaptive responses to both routine and emergency situations.
III-B Reward for Platooning
The reward function is central to shaping the behavior of autonomous vehicle platoons. Inspired by the social behavior in AVs[26], a reward system is designed to account for both individual vehicle performance and the collective dynamics of the platoon as a whole. This approach ensures that each vehicle operates safely and efficiently while maintaining coordination and cohesion within the platoon.
The total reward $R_{global}(s,a)$ is formulated as the sum of individual rewards and system-wide collaborative rewards:
$R_{global}(s,a)=\frac{1}{n}\sum_{i=1}^{n}R_{ind}^{i}+R_{sys}$ (8)
Both the individual vehicle perspective and the platoon-wide perspective are considered in the reward design. From the individual vehicle’s perspective, the reward function emphasizes safety and efficiency, focusing on collision avoidance, lane keeping, speed control, and smooth acceleration: